OpenStudy (mathmusician):

Can someone tell me the range of this function? I don't know how it has been too long

OpenStudy (mathmusician):

\[\cos^{-1} (\frac{ sinx }{ \sqrt{2} })\]

OpenStudy (mathmusician):

@ShadowLegendX